(1 point) Solve the separable differential equation for u e5u+8t Use the following initial condition: u(0) = 15. U =
Answer #1

esu. sutsturo)=15 du = e 54t8t e54 e8t dt du = e su est dt Je-sudu = sestat integrating both side wor.to u and t respectivelye-su = taking sest + 8 + 0-75 in both sides we get - 5 u = ln (-sest + 5+ 8 0-75 1 u = - = ln(-5e8t +5+ 86-75) u = - en -sest